OpenStudy (anonymous):

evaluate: sec^2(-13pie/6)-csc^25pie/4

OpenStudy (anonymous):

\[\sec ^2(-13\pi/6)=\sec ^2(\pi/6)\] \[\csc ^2(5\pi/4)=\sec ^2(\pi/4)\] now put corresponding values i.e (4/3)-2